# Hands-on STEM Learning Experiences using Digital Technologies

### v1

@@ III.3 DIY Lift Generator for STEM: Tubora
The versatility of this exhibit is noteworthy. It allows children in the first years of primary school (aged 6-8) to construct paper objects in a creative manner, thereby determining the probability of them flying out of the tube. This provides students with the opportunity to experiment with different shapes, testing them each time, and to color them in an imaginative way. In addition to cones and paper airplanes, students can investigate the efficacy of twirling paper helicopters. This involves varying factors such as the type of paper, the addition of paperclips, the size of wings, and so forth.
Figure 4: STEM Engineering Activity: the runaway zoo. Examples of activity sheets.
@@
they are presented with an object (specifically a small lightweight stuffed animal) that is too heavy to be lifted by the fan alone. Each group is provided with the same materials (paper, tulle, cotton and umbrella fabric, plastic bags, string, masking tape) and must design the most efficient solution to allow the “skydiver” to rise and escape from the tube. The students will have the opportunity to experiment with different approaches to the design of the parachute canopy. They will select the material they wish to use for the main canopy and then proceed through the engineering design process, which involves defining the problem, brainstorming potential solutions, creating prototypes and testing them in a wind tunnel. Each group will also be required to complete a record sheet, which will help them document their choices, the reasons behind them, and some background information on the plush animal and its need to escape. They are also required to draft sketches of the front and top views of their prototype to complete their task thus fulfilling the final step of the Engineering design process (document and share) –see Fig.4. This parachute challenge offers the opportunity to investigate a multitude of aspects related to flight and the effects of force on an object. It also provides an opportunity to examine the engineering design process, which is a systematic approach to problem-solving.
III.4 Highly Visual Cloud Chamber –Physics in Action!
A facility for studying tiny invisible elementary particles that enter our atmosphere from the outer space (cosmic rays) or that are generated by the decay of radioactive elements within the Earth itself, and then pass right through our bodies, is the well-known apparatus known as “Wilson’s cloud chamber”. This apparatus was invented in the early 1900s by Charles Wilson as one of the first particle detectors. The majority of these particles are generated by the decay of natural radon found in the air (itself a product of radioactive element’s decay in the ground) while a certain percentage, which varies with altitude, does originate from the cosmic radiation we receive from the outer space (primary and secondary cosmic rays).
Figure 5: Typical particles tracks in a cloud chamber device (Left). Full video available at https://www.youtube.com/watch?v=IwyLzqrRJbw Right: Public interest on the DIY cloud chamber.
To build a cloud chamber for educational purposes one needs a transparent box with a black bottom metal tray maintained at a temperature of approximately minus 30 degrees Celsius, accompanied by the presence of a stable cloud of minute droplets of isopropyl alcohol, typically generated by condensation from a liquid state. After a period of time, the top lid of the chamber, which has been soaked in warm alcohol, produces a cloud that reaches the cold metal bottom. When a charged particle passes through at a high speed it generates a visible track of larger droplets by a mechanism known as coalescence. The droplets of alcohol can then be detected due to their ability to diffuse the light emitted by the strips of LEDs positioned on the sides, as illustrated in Fig.5.
The particles that can be observed in a cloud chamber can be broadly classified into three categories: alpha particles, beta particles and muons. Alpha particles are composed of two protons and two neutrons. They have a short, fat track and are produced by the decay of radon. Beta particles are electrons (or anti-electrons, a.k.a. positrons), usually following long straight paths or zigzag paths, depending on the energy and momentum of the particles. Muons are particles emitted from the decay of secondary cosmic rays and can be observed as short vertical tracks. Further details on the construction and operation of a cloud chamber can be found in [9].
III.5 Galileo and Leonardo Mechanical Clocks
The first application of a pendulum in a clock used an escapement system is attributed to Galileo Galilei from his original drawings. Following the discovery of the isochronism of a pendulum by Galileo’s observation of a swinging lamp in Pisa Cathedral (dating back to 1600), a working gravity-powered clock can be made on a laser cutter with the addition of standard bearings and a few parts made from plywood. As shown in Fig.6 (Left), the design of a Galileo mechanical clock can be written using OpenSCAD –a free, fully customizable software.
Figure 6: OpenSCAD design of a Galileo mechanical clock (Left). Video demo available at: https://www.youtube.com/watch?v=a6B1bxgxAc4 Right: Leonardo Mechanical Clock. Demo video available at: https://www.youtube.com/shorts/g4hkIHu89kE
Galileo’s clock mechanism uses gravitational potential energy as its power source, and carries two arcs; a small lower arc that receives the impulse, while an upper wide arc lifts a detent. This model has been designed in collaboration with the “Amici dell’orologeria Pesarina” association, engaged since many years in the dissemination of knowledge concerning watchmaking, a tradition present in the area of Pesariis (a ancient village in the mountains of Friuli Venezia Giulia region, Italy) since 1600, and its dissemination in local schools. The request was to produce a model escapement that could be made cheaply and then to introduce a kit into secondary schools so that students could experience the mechanisms first hand. The files for this build can be dowloaded freely (https://www.thingiverse.com/thing:5549040) and can be further customized in OpenSCAD easily.
The Leonardo Da Vinci Clock depicted in Fig.6 (Right) is another wooden pendulum clock that can be constructed in a FabLab for educational purposes (https://www.thingiverse.com/thing:4328105). This clock model is powered by a horizontal rotary pendulum, and utilizes the weight contained in a canister and some gears to measure time. The canister has stored potential energy gained by raising it up against gravity. The potential energy is converted to kinetic energy allowing the gears to move their parts. By adjusting the weight and balance of the pendulum, the clock’s speed can then be controlled.
III.6 Giant Galton Board
A large Galton board allows students to visualize the preferred trajectories of falling (ping-pong) balls while being fed into the device from the top at arbitrary rates. Due to the differing initial conditions established by the participants, the gravitational force and the presence of ordered, fixed obstacles along their paths, the individual balls experience random elastic collisions as they descend. The falling balls positions will not concentrate at specific locations: Instead, they will be characterized by a distribution (with one or more peaks) resulting from the influence of the stochastic effects due to the presence, or absence, of the obstacles. The spatial distribution of the falling balls with homogeneous obstacles is observed to exhibit a symmetric bell-shaped (Gaussian) curve –see Fig.7.
Figure 7: Giant Galton board of descent ping-pong balls.
The observation of individual ball trajectories represents the dynamics of a single ensemble member, whereas the bell statistical distribution of the final balls positions represents the total ensemble average for the system. This device can serve as a learning platform for explaining weather forecast models, as they use the equations governing the atmosphere in conjunction with a stochastic algorithm called “Montecarlo” to generate a weather forecast (corresponding to the resulting distribution of balls). The model, as well as the device, exhibits a high degree of sensitivity to its initial conditions (of the dropped balls), where even minor initial perturbations give rise to divergent weather predictions. The presence or absence of the obstacles correspond to the various parameters of the underlying model, and they can be moved to “tune” the results.
III.7 Augmented Reality Sandbox
The Augmented Reality (AR) Sandbox is an educational hands-on exhibit developed by the University of California, Davis, USA. It combines a box filled with real sand, with virtual topography and simulated water. The apparatus is constructed using a Microsoft Kinect 3D camera, in addition to a video projector and a Linux PC to achieve extensive real-time computer simulations and visualization. The AR Sandbox allows the reconstruction of geographical models through the shaping of the sand. These models are then augmented by an elevation color map, topographic contour lines, and simulated liquid water.
The device, illustrated in Fig.8, has been employed for science communication projects involving experiments on geological hazards and volcano alerts. The system is useful for illustrating concepts related to modeling volcanoes, including the processes involved in volcano deformation and the study of magma reservoirs.
A variety of hands-on experiments can be conducted, such as placing a small, inflated balloon beneath the sand. Deflating the balloon can then demonstrate the effect of a collapsing volcano magmatic chamber. It is also possible to illustrate the impact of the rising sea level on coastal areas: for instance Easter Island –located in the southeastern Pacific Ocean, is highly vulnerable to the effects of climate change, and this can be visually demonstrated with great clarity by varying digitally the height of the sea level.
III.8 2D-to-3D Image Conversion via Artificial Intelligence Model
3D visualization has the potential to enhance the understanding of STEM among young students, thereby facilitating more effective learning. Virtual learning opens up new worlds in STEM education, particularly in the context of abstract concepts and the development of specific experiences [10]. Virtual learning can facilitate the exploration of the interior of a human cell, traverse planetary surfaces, or navigate subterranean magma chambers. These are examples of the ”intangible” world that can be experienced in virtual learning environments. In order for this to be achieved, it is necessary to recreate a realistic 3D environment. It would be even more beneficial if the viewer were able to perceive this content more naturally, i.e., without the need for special 3D glasses.
Figure 9: Outputs of ”altiro3D” algorithm to represent reality. Left: original single RGB image. Center: multiview synthesis of a single image carried out through artificial intelligence. Right: 3D display on a headset-free Lenticular 3D screen. Video demo available at: https://www.youtube.com/shorts/hJDVb2TzBr0
A visual 3D scene representation of a learning environment serves as a medium through which the integration of STEM learning may be effectively transformed into a tangible and meaningful experience. It has a tremendous potential to revolutionize education, yet it remains to be further explored. The free library ”altiro3D, which enables the representation of reality from a single 2D RGB image or flat video, has also been developed in our Scientific FabLab [11]. This software allows the synthesis of multiple N-viewpoints images or video to create a realistic (real-time) 3D immersive experience when displayed on a lenticular, free-view LCD screen as shown in Fig.9. altiro3D does not require heavy computing runtime and can support a wide range of application scenarios in education and science. The synthesis of multiple views from a single image is achieved through the integration of machine learning together with deep learning algorithms, and simple inpainting techniques, which map all image pixels.
IV Discussion
@@
A further challenge is posed by the increasing number of students from other countries: in the 2021/2022 academic year, 10.6% of students in Italy were from other countries, with the figure in Friuli Venezia Giulia (FVG) Region reaching 15.4% and 22.7% in the province of Gorizia [13]. A considerable number of educational institutions are currently undergoing renewal works, with classes being temporarily relocated to smaller classrooms or to alternative buildings that have been converted into schools. The challenges of working in such circumstances are considerable, particularly when the classes are overcrowded and a significant proportion of students lack proficiency in the local language. STEM curricula are highly engaging for students, particularly those with language difficulties. The hands-on nature of many STEM subjects enables all students to participate fully, regardless of their linguistic abilities. This phenomenon is further evidenced when students with disabilities are present in the classroom. According to our experiences, STEM activities can be modified to accommodate the specific needs of these students, with the objective of overcoming physical limitations and addressing mental impairments. During a lesson in which a class with a blind student was being taught, the possibility of combining a buzzer with the LED outcome of the experiment proved invaluable in fully involving the student in the lesson.
On another occasion, a student in a wheelchair was able to participate completely in the activity by having all the educational material relocated to the correct height for him to reach. In another instance, a student with an autism spectrum disorder had the time needed to fully appreciate the effects of air thrust on a small parachute, while his fellow students were engaged in the collection of data and the preparation of a subsequent demonstration.
@@
The experiences of one of the authors (G.F.) on providing one-off, hands-on learning activities in primary and secondary schools over the past decade has been predominantly focused on the delivery of engaging, 1.5-3 hour long sessions that engage the entire class, while also providing all the necessary materials. This approach has the dual benefit of relieving teachers of the burden of providing materials while also allowing for more focused and effective learning.
@@
Figure 10: Students STEM activities for Italian Schools from the FVG Region carried out at the Scientific FabLab of ICTP.
@@
Students currently engaged in primary and secondary education have experienced the challenges of the Covid-19 lockdown and a prolonged period of distance learning, which has resulted in significant difficulties in working in groups and reduced dexterity. This kind of learning hands-on experience using digital technologies, can be extremely useful in addressing both these problems. In our experience, dividing the class into four groups is optimal, as it allows for more focused and effective learning while allowing a single educator (such as the situation usually presents) to monitor the progress of each group, ensuring that all groups remain on track while providing additional challenges to the more advanced group.
@@
While working with smaller groups or a one-to-one approach would also be interesting, the limitations posed by the need to engage the whole class and have a single educator limit the possibilities, resulting in an ideal configuration with four groups of students, forcing them to work as teams, exchanging ideas and comparing their different approaches, allowing for a more focused and effective learning. Furthermore, this configuration has enabled teachers to examine the interactions between their students in a group setting, offering valuable insights. The decision regarding the composition of groups is always a challenging one, with students frequently attempting to maintain proximity to their acquaintances, thereby creating groups that are not optimally balanced. When working with primary school pupils the involvement of teachers, who possess a comprehensive understanding of the students and are capable of dividing them (maintaining the separation of more challenging students and creating more balanced groups) can be invaluable in this regard. In the context of working with secondary school students, it is possible to allow them to choose their own group, thereby giving them the opportunity to assume this responsibility.
@@
In most cases, the curricula teachers participate, assisting students with more difficulties and providing important context with references to lessons already given on similar topics. Frequently, they choose to discuss in advance the topics covered during such activities, stimulating students’ curiosity before the laboratory, creating an optimal learning environment. The class is often requested to provide feedback on the STEM topics presented, which may take the form of a simple drawing or collaborative poster for primary school students, or a summary or more creative medium such as a video, comic, rap song, newscast or other form for secondary students. The variety of media employed and the creativity displayed by classes that have previously encountered disciplinary problems during the laboratory is often noteworthy.
V Conclusions
@@
STEM students can greatly benefit when receiving training and education based on tangible and visual learning. The prototypes presented in this work as examples allows to put concrete experiences in their hands. The aforementioned examples have been generated utilizing digital technologies that are readily accessible in any FabLab worldwide. We have discussed and assessed some of our experiences and activities done in Italian Schools using such devices. For example, by touching complex 3D structures, visualizing cosmic rays, and building the needed blocks to learn micro-controllers in class.
@@
We have verified how hands-on STEM learning emphasizes students potential and opens a new path to present them with a smooth connection between the understanding of abstract scientific phenomena and a deeper, more comprehensive learning experience. This novel approach in action is appealing to scholars and can sustain new generations of scholars in an inclusive way facilitating the quality of STEM education for all.
Acknowledgement
@@
We thank Sara Sossi, Erika Ronchin and Marco Baruzzo who have helped with our ICTP SciFabLab STEM activities in the last few years. Sincere gratitude is also due to Fondazione Pietro Pittini (FPP) for partnering with us in many of these projects.

### v2

@@ 4.2 DIY lift generator for STEM: Tubora
The versatility of this exhibit is noteworthy. It allows children in the first years of primary school (aged 6–8) to construct paper objects in a creative manner, thereby determining the probability of them flying out of the tube. This provides students with the opportunity to experiment with different shapes, testing them each time, and to color them in an imaginative way. In addition to cones and paper airplanes, students can investigate the efficacy of twirling paper helicopters. This involves varying factors such as the type of paper, the addition of paperclips, the size of the wings, and so forth.
Figure 9: Examples of activity sheets for the STEM Engineering activity. The Runaway Zoo.
@@
They are presented with an object (specifically a small, lightweight stuffed animal) that is too heavy to be lifted by the fan alone. Each group is provided with the same materials (paper, tulle, cotton and umbrella fabric, plastic bags, string, masking tape) and must design the most efficient solution to allow the “skydiver” to rise and escape from the tube. The students will have the opportunity to experiment with different approaches to the design of the parachute canopy. They will select the material they wish to use for the main canopy and then proceed through the engineering design process, which involves defining the problem, brainstorming potential solutions, creating prototypes and testing them in a wind tunnel. Each group will also be required to complete a record sheet, which will help them document their choices, the reasons behind them, and some background information on the plush animal and its need to escape. They are also required to draft sketches of the front and top views of their prototype to complete their task thus fulfilling the final step of the engineering design process (document and share), see Figure 9. This parachute challenge offers the opportunity to investigate a multitude of aspects related to flight and the effects of force on an object. It also provides an opportunity to examine the engineering design process, which is a systematic approach to problem-solving.
5 Discussion
One of the objectives of this work is to include an account of some of our experiences and activities carried out in Italian schools using digital technologies from FabLabs. Our school experiences and public activities are not only directed to school-aged childrens, but are also being developed for interested scholars of all ages.
@@ 5 Discussion
A further challenge is posed by the increasing number of students from other countries, in the 2021/2022 academic year, 10.6% of students in Italy were from other countries, with the figure in the FVG region reaching 15.4% and 22.7% in the province of Gorizia [13]. A considerable number of educational institutions are currently undergoing renewal work, with classes being temporarily relocated to smaller classrooms or to alternative buildings that have been converted into schools. The challenges of working in such circumstances are considerable, particularly when the classes are overcrowded and a significant proportion of students lack proficiency in the local language. The ages of participants in our activities during the academic year 2023/24 were as follows. As far as laboratory hand-ons classes were concerned, there were 172 workshops carried out for the primary school years 1–5 (students having 6 to 11 years of age in correspondence). The age groups consisted of 12% in the 1st class, 15% in the 2nd, 19% in the 3rd, 24% in the 4th, 24% in the 5th year, and 6% in multiple age groups classes.
STEM curricula are highly engaging for students, particularly those with language difficulties. The hands-on nature of many STEM subjects enables all students to participate fully, regardless of their linguistic abilities. This phenomenon is further evidenced when students with disabilities are present in the classroom. According to our experiences, STEM activities can be modified to accommodate the specific needs of these students, with the objective of overcoming physical limitations and addressing mental impairments. During a lesson in which a class with a blind student was being taught, the possibility of combining a buzzer with the LED outcome of the experiment proved invaluable in fully involving the student in the lesson.
On another occasion, a student in a wheelchair was able to participate completely in the activity by having all the educational materials relocated to the correct height for him to reach. In another instance, a student with an autism spectrum disorder had the time needed to fully appreciate the effects of air thrust on a small parachute, while his fellow students were engaged in the collection of data and the preparation of a subsequent demonstration.
@@
The experiences of one of the authors (G.F., also an assistant in the SciFabLab) on providing one-off, hands-on learning activities in primary and secondary schools over the past decade has been predominantly focused on the delivery of engaging, 1.5 to 3-hour-long sessions that engage the entire class, while also providing all the necessary materials. This approach has the dual benefit of relieving teachers of the burden of providing materials while also allowing for more focused and effective learning.
@@
Figure 10: Students’ STEM activities for Italian schools from the FVG region carried out at the Scientific FabLab of ICTP.
@@
Students currently engaged in primary and secondary education have experienced the challenges of the Covid-19 lockdown and a prolonged period of distance learning, which has resulted in significant difficulties in working in groups and reduced dexterity. This kind of learning, utilizing hands-on experience using digital technologies, can be extremely useful in addressing both of these problems. In our experience, dividing the class into four groups is optimal, as it allows for more focused and effective learning while allowing a single educator (which is the typical situation) to monitor the progress of each group, ensuring that all groups remain on track while providing additional challenges to the more advanced group.
@@
While working with smaller groups or a one-to-one approach would also be interesting, the limitations posed by the need to engage the whole class with a single educator limit the possibilities. An ideal configuration would have four groups of students, forcing them to work as teams to exchange ideas and compare their different approaches, allowing for a more focused and effective learning environment. Furthermore, this configuration has enabled teachers to examine the interactions between their students in a group setting, offering valuable insights. The decision regarding the composition of groups is always a challenging one, with students frequently attempting to maintain proximity to their acquaintances, thereby creating groups that are not optimally balanced. When working with primary school pupils, the involvement of teachers, who possess a comprehensive understanding of the students and are capable of dividing them (maintaining the separation of more challenging students and creating more balanced groups), can be invaluable in this regard. In the context of working with secondary school students, it is possible to allow them to choose their own groups, thereby giving them the opportunity to assume this responsibility.
@@
In most cases, the curricula teachers participate, assisting students that are having difficulty and providing important context with references to lessons already given on similar topics. Frequently, they choose to discuss in advance the topics covered during such activities, stimulating students’ curiosity before the laboratory, creating an optimal learning environment. The class is often requested to provide feedback on the STEM topics presented, which may take the form of a simple drawing or collaborative poster for primary school students, or a summary or more creative medium such as a video, comic, rap song, newscast, or other form for secondary students. The variety of media employed and the creativity displayed by classes that have previously encountered disciplinary problems during the laboratory is often noteworthy.
A recent extensive literature review indicates that digital technology can positively affect STEM education in terms of knowledge or skill acquisition and learning engagement in young children [14]. This fact was found regardless of gender but highly relevant to age. Learning from, with, and through technology occurs when the technology is easy to be used as a tool to deliver knowledge. In most cases, these technologies become limited within the context of early childhood education since they include computer programming, simulation software, robotics, 3D technology, video games, apps., etc. The new trends of emerging digital technologies, such as AI, in the STEM learning context become limited in children’s education [15]. An important problem outlined in these existing approaches is the age of students. Our study differentiates with such reports because we actually motivate the primary and secondary schools to participate in STEM festivals. Ours are preliminary outcomes based on our experiences not only carried out in Italian schools but also during large, free public activities like Maker Faires and Science Picnics which allow anyone, of any age, to explore scientific concepts in simple, innovative ways. By this, we encourage the young science scholars to be creative and incentivize them to develop their own abilities for research, discovery and ingenuity.
Our goal in this article is to introduce readers on the importance of adopting digitally controlled technologies found in FabLabs for tangible and visual STEM learning. Educators may adjust their standard practices to incorporate hands-on STEM learning activities into their curriculum at their own pace and own needs. Not all scientific prototypes illustrated are for school-aged children, as with the Cloud Chamber and the AI-based 3D imaging. We aim at getting better understanding of science through the utilization of tangible and visual examples as those illustrated and explained in our paper. In our study we keep in mind the relevant work carried out by teachers as well as aim to reach a larger range of professionals, decision-makers, politicians, scientists and interested students in general. This works aims to be of benefit for the progress of society as a whole.
6 Conclusions
@@
STEM students can greatly benefit when receiving training and education based on tangible and visual learning. The prototypes presented in this work as examples allows us to put concrete experiences in their hands. The aforementioned examples have been generated utilizing digital technologies that are readily accessible in any FabLab worldwide. We have discussed and assessed some of our experiences and activities done in Italian schools and large public events using and demonstrating such devices such as touching complex 3D structures, visualizing cosmic rays, and building the needed blocks to learn microcontrollers in class.
At present, it is difficult to compare the similarities and differences between our research on tangible and visual STEM learning experiences and existing studies since there are not many similar studies. It is premature to derive definitive conclusions because more data within different academic environments and experiences around the world are needed.
@@
We have built upon and verified how visual and tangible STEM learning emphasizes student potential and opens a smooth connection between the understanding of abstract scientific phenomena and a deeper, more comprehensive learning experience. This novel approach in action is appealing to scholars and can sustain new generations of young science scholars in an inclusive way facilitating the quality of STEM education for all.
Author contributions
Gaia Fior: Conceptualization, Methodology creation, Data Curation, Validation, Writing - review & editing. Carlo Fonda: Conceptualization, Methodology creation, Investigation, Resources, Writing - review & editing. Enrique Canessa: Conceptualization, Methodology creation, Investigation, Formal analysis, Writing - review & editing. Supervision. All authors have approved the final version of the manuscript for publication.
Use of Generative-AI tools declaration
The authors declare they have not used Artificial Intelligence (AI) tools in the creation of this article.
Acknowledgments
@@
The authors would like to thank Sara Sossi, Erika Ronchin, and Marco Baruzzo who have helped with their ICTP SciFabLab STEM activities during the last few years. Sincere gratitude is also due to Fondazione Pietro Pittini (FPP) for partnering with us in many of these projects.
Conflict of interest
The authors declare no conflict of interest in this paper.